OpenStudy (anonymous):

The length of a rectangle is 3 times its width. If the length were increased by 2 inches and the width by 1 inch, the new perimeter would be 62 inches. What is the width and length of the rectangle? If w represents the width of the rectangle, then which equation could be used to solve for w? A.(w + 1)(3w + 2)=62 B.2(4w + 3) = 62 C.2(3w + 3) = 62

OpenStudy (campbell_st):

well let the width be w then the length is 3w add 2 to the length so l = 3w + 2 add 1 to the width so w = w + 1 now perimeter is P = 2w + 2l so substituting the known information 62 = 2(w + 1) + 2(3w +2)

OpenStudy (campbell_st):

or using P = 2(l + w) 62 = 2(3w + 2 + w + 1) this maybe easier to simplify